When temperature of a gas is 20°C and pressure is changed from P1 = 1.01*105 Pa to P2 =1.165 * 105 Pa then the volume changed by 10%.The bulk modulus is
        (a) 1.55 * 105 Pa
        (b) 0.115 * 105 Pa
        (c) 1.4 * 105 Pa
        (d) 1.01 * 105 Pa

Grade:12th pass

1 Answers

Aditi Chauhan
9 years ago

Sol :- (a) B= ?p/?V/V = (1.165 * 105 – 1.01*105)/0.1

= 1.55 * 105 Pa